OpenStudy (anonymous):

A circular jogging track is being designed so that one complete lap around the track is one mile long. What should the radius of the track be to the nearest foot? (one mile equals five thousand two hundred eighty feet) two thousand six hundred forty feet one thousand seven hundred sixty feet one thousand six hundred eighty-two feet eight hundred forty-one feet

OpenStudy (anonymous):

i need help asap

OpenStudy (anonymous):

i need help asap

OpenStudy (ashleyisakitty):

r=(C/2*pi) r=[5280/(2*3.14)] r= ?

OpenStudy (anonymous):

wut

OpenStudy (anonymous):

can you tell me the answer

OpenStudy (ashleyisakitty):

What does the last equation equal to?

OpenStudy (anonymous):

i get you a medal if you answer it

OpenStudy (ashleyisakitty):

I did.

OpenStudy (ashleyisakitty):

ill give you a medal if you use your brain

OpenStudy (anonymous):

tell me it

OpenStudy (anonymous):

mean

OpenStudy (ashleyisakitty):

[5280/(2*3.14)] equals what

OpenStudy (anonymous):

two thousand six hundred forty feet

OpenStudy (ashleyisakitty):

uh, no.

OpenStudy (ashleyisakitty):

do you know of those magical things called calculators?

OpenStudy (anonymous):

oh let me try agan

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

im 12

OpenStudy (ashleyisakitty):

your age doesnt give you an excuse for not using common sense

OpenStudy (anonymous):

YES i know

OpenStudy (anonymous):

is it 2 times 3 times 14

OpenStudy (ashleyisakitty):

3.14 is pi

OpenStudy (ashleyisakitty):

so its 2 x 3.14

OpenStudy (anonymous):

so you 2x3

OpenStudy (anonymous):

=6

OpenStudy (ashleyisakitty):

no, 2 * 3.14, do it exactly.

OpenStudy (anonymous):

whhat does * mean

OpenStudy (ashleyisakitty):

Its multiplication. Multiply 2 by 3.14 exactly, not just 2 by 3.

OpenStudy (anonymous):

can you just expplann it be working it out beacuse i do not get it

OpenStudy (ashleyisakitty):

Do you not know basic multiplication, sweetheart? its just 2 by 3.14. 3.14 2 _____ 6.28

OpenStudy (anonymous):

I DO

OpenStudy (ashleyisakitty):

Now that you have that, divide 6.28 into 5280.

OpenStudy (anonymous):

and when i get that will it be my answer

OpenStudy (ashleyisakitty):

what do you get when you divide it?

OpenStudy (anonymous):

840

OpenStudy (ashleyisakitty):

Yes, correct. That is your answer.

OpenStudy (anonymous):

thank you sooo much

OpenStudy (ashleyisakitty):

yerr welcome